# Noise-Tolerant Hybrid Approach for Data-Driven Predictive Control

### v1

@@ 4.1 Noise-Free Case
We analyze the stability of NTDPC in the noise-free case ($\mathbf{\zeta}_{\text{ini}}(k)=0$), tracking references $(r_{y},r_{u})$.
@@
There exist $\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ such that for all $(y,u)\in\mathcal{Y}_{i}\times\mathcal{U}_{i}$, where $\mathcal{Y}_{i}\subseteq\mathbb{R}^{n_{y}}$ and $\mathcal{U}_{i}\subseteq\mathbb{R}^{n_{u}}$ are the projections of $\mathcal{Y}$ and $\mathcal{U}$ onto the $i$-th prediction step:
$\alpha_{1}(\|\text{col}(y-r_{y},u-r_{u})\|)\leq l(y,u)\leq\alpha_{2}(\|\text{col}(y-r_{y},u-r_{u})\|),$
where $l(y,u)=\|y-r_{y}\|_{Q_{i}}^{2}+\|u-r_{u}\|_{R_{i}}^{2}$, and $(r_{y},r_{u})\in\mathcal{Y}_{i}\times\mathcal{U}_{i}$ is a constant admissible equilibrium.
@@
For any admissible initial state $\mathbf{z}_{\text{ini}}(k)\in\mathcal{U}^{T_{\text{ini}}}\times\mathcal{Y}^{T_{\text{ini}}}$, there exist $\rho\in\mathcal{K}_{\infty}$ with $\rho<\text{id}$, a prediction horizon $N\geq T_{\text{ini}}$, and $\hat{u}(N|k)\in\mathcal{U}_{N}\subseteq\mathbb{R}^{n_{u}}$ such that $\hat{y}(N|k)\in\mathcal{Y}_{N}\subseteq\mathbb{R}^{n_{y}}$:
$l(\hat{y}(N|k),\hat{u}(N|k))\leq(\text{id}-\rho)\circ l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})).$
@@
Problem 4 is feasible for all $k\geq T_{\text{ini}}$ with $\mathbf{\zeta}_{\text{ini}}(k)=0$, and the optimal slack satisfies $\sigma_{y}^{*}(k)=0$. The sets $\mathcal{U},\mathcal{Y}$ contain the equilibrium $(r_{y},r_{u})$.
@@
Under Assumptions 4.1, 4.2, and 4.3, with $\mathbf{\zeta}_{\text{ini}}(k)=0$ and a constant admissible equilibrium $(r_{y},r_{u})$, the closed-loop system under NTDPC is asymptotically stable, i.e., $y(k)\to r_{y}$, $u(k)\to r_{u}$ as $k\to\infty$.
@@
Define the non-minimal state as in Definition 4.2:
$\mathbf{z}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}}),\ldots,y(k-1),u(k-T_{\text{ini}}),\ldots,u(k-1)).$
The error state is $\tilde{X}_{\text{ini}}(k)=\mathbf{z}_{\text{ini}}(k)-\text{col}(\mathbf{r}_{y},\mathbf{r}_{u}),$ where $\mathbf{r}_{y}=\text{col}(r_{y},\ldots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\ldots,r_{u})$. Define the auxiliary function as $W(\mathbf{z}_{\text{ini}}(k))=\sum_{i=1}^{T_{\text{ini}}}l(y(k-i),u(k-i)),$ where the stage cost is $l(y(i|k),u(i|k))=\|y(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u(i|k)-r_{u}\|_{R_{i}}^{2},$ with $Q_{i},R_{i}\succ 0$. The optimal cost of Problem 4 is $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k)),$ since $\sigma_{y}^{*}(k)=0$ in the noise-free case (Assumption 4.3). Consider the storage function [24]:
$V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k)).$
To show positive definiteness of $V(\mathbf{z}_{\text{ini}}(k))$, using Assumption 4.1, we have:
$l(y(k-i),u(k-i))\geq\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$
@@
so $W(\mathbf{z}_{\text{ini}}(k))\geq\sum_{i=1}^{T_{\text{ini}}}\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|).$ Since $\tilde{X}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}})-r_{y},\ldots,u(k-1)-r_{u})$, there exists $\alpha_{1,W}\in\mathcal{K}_{\infty}$ such that $W(\mathbf{z}_{\text{ini}}(k))\geq\alpha_{1,W}(\|\tilde{X}_{\text{ini}}(k)\|).$ Since $J^{*}(k)\geq 0$, there exists $\alpha_{1,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\geq\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ For the upper bound:
$l(y(k-i),u(k-i))\leq\alpha_{2}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$
so $W(\mathbf{z}_{\text{ini}}(k))\leq T_{\text{ini}}\alpha_{2}\left(\max_{i=1,\dots,T_{\text{ini}}}\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|\right).$ For $J^{*}(k)$, the predictor output $\hat{\mathbf{y}}_{N}^{*}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini}}(k)+\mathbf{P}_{2}\mathbf{u}_{N}^{*}(k)$ (Lemma 3.5) depends on $\mathbf{z}_{\text{ini}}(k)=\tilde{X}_{\text{ini}}(k)+\text{col}(\mathbf{r}_{y},\mathbf{r}_{u})$. Bounded sets $\mathcal{Y},~{}\mathcal{U}$ (Assumption 4.3) and the predictor’s linearity ensure that $\|y^{*}(i|k)-r_{y}\|$, $\|u^{*}(i|k)-r_{u}\|$ are bounded by $\|\tilde{X}_{\text{ini}}(k)\|$. Thus, there exists $\alpha_{2,J}\in\mathcal{K}_{\infty}$ such that $J^{*}(k)\leq\alpha_{2,J}(\|\tilde{X}_{\text{ini}}(k)\|).$ Combining, there exists $\alpha_{2,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ Hence, $\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|)\leq V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ Next, by defining the supply function as:
$s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))=l(\hat{y}(N|k),\hat{u}(N|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})),$
where $\hat{y}(N|k)$, $\hat{u}(N|k)$ satisfy Assumption 4.2. Construct a suboptimal solution at $k+1$:
$\mathbf{u}_{s}(k+1)$ $=\text{col}(u^{*}(1|k),\ldots,u^{*}(N-1|k),\hat{u}(N|k))$
$\mathbf{y}_{s}(k+1)$ $=\text{col}(y^{*}(1|k),\ldots,y^{*}(N-1|k),\hat{y}(N|k)),$ (53)
with $\sigma_{y,s}(k+1)=0$. The suboptimal cost is as $J_{s}(k+1)=\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k)).$ Since $J^{*}(k)=\sum_{j=0}^{N-1}l(y^{*}(j|k),u^{*}(j|k))$ and $J^{*}(k+1)\leq J_{s}(k+1)$:
$J^{*}(k+1)\leq J^{*}(k)-l(y^{*}(0|k),u^{*}(0|k))+l(\hat{y}(N|k),\hat{u}(N|k))$ (54)
For the auxiliary function $W$:
$W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))=l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (55)
Using (61) and (55), thus:
$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))=[J^{*}(k+1)-J^{*}(k)]+[W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))]$ (56)
Then, we can rewrite the preceding equality in (56) as:
$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))$ $\leq[J^{*}(k)-l(y^{*}(0|k),u^{*}(0|k))+l(\hat{y}(N|k),\hat{u}(N|k))]$
$\quad-J^{*}(k)+[l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))]$
$=s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (57)
By Assumption 4.2, $s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))\leq-\rho\circ l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})).$ Since $l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))\geq\alpha_{1}(\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|)$, we can rewrite the inequality (57) as:
$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))\leq-\rho\circ\alpha_{1}(\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|)\leq-\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|).$
Since $V$ is positive definite and radially unbounded with respect to $\tilde{X}_{\text{ini}}(k)$, and satisfies a strict dissipation inequality, standard Lyapunov results [25] imply $\tilde{X}_{\text{ini}}(k)\to 0$ as $k\to\infty$. Thus, $y(k)\to r_{y}$, $u(k)\to r_{u}$, proving asymptotic stability. Since $V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)$ and $W(\mathbf{z}_{\text{ini}}(k))\geq 0$:
$J^{*}(k)\leq V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|),$
ensuring the cost is bounded by a $\mathcal{K}_{\infty}$ function of the error state. This completes the proof. ∎
@@
The bound $J^{*}(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)$ satisfies the requirement for a $\mathcal{K}_{\infty}$ cost bound, derived within the storage function analysis.
4.2 Noisy Case: Robustness to Measurement Noise
We analyze the stability of NTDPC under measurement noise ($\mathbf{\zeta}_{\text{ini}}(k)\neq 0$), tracking constant admissible references $(r_{y},r_{u})$. The system is shown to be input-to-state stable (ISS) with respect to $\mathbf{\zeta}_{\text{ini}}(k)$, following the framework in [2, 25].
Assumption 4.5 (Bounded Noise Variance).
There exists $\bar{c}>0$ such that $\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}]\leq\bar{c}$ for all $k$, where $\mathcal{F}_{k}$ is the filtration up to time $k$.
Assumption 4.6 (Lipschitz Predictor).
There exists $L_{P}>0$ such that:
$\|\mathbf{P}_{1}(\mathbf{z}_{\text{ini},m}(k)-\mathbf{z}_{\text{ini}}(k))\|\leq L_{P}\|\mathbf{\zeta}_{\text{ini}}(k)\|,$
where $L_{P}=\|\mathbf{P}_{1}\|_{2}$ is the spectral norm of $\mathbf{P}_{1}$, and $\mathbf{z}_{\text{ini},m}(k)=\mathbf{z}_{\text{ini}}(k)+\text{col}(0,\mathbf{\zeta}_{\text{ini}}(k))$.
Assumption 4.7 (Terminal Stabilizing Condition - Noisy).
For any admissible initial state $\mathbf{z}_{\text{ini}}(k)\in\mathcal{U}^{T_{\text{ini}}}\times\mathcal{Y}^{T_{\text{ini}}}$, there exist $\rho\in\mathcal{K}_{\infty}$ with $\rho<id$, a prediction horizon $N\geq T_{\text{ini}}$, and $\hat{u}(N|k)\in\mathcal{U}_{N}\subseteq\mathbb{R}^{n_{u}}$ such that $\hat{y}(N|k)\in\mathcal{Y}_{N}\subseteq\mathbb{R}^{n_{y}}$:
$\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]\leq(\text{id}-\rho)\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$
Assumption 4.8 (Feasibility - Noisy).
Problem 4 is feasible with probability at least $1-\epsilon$, for small $\epsilon>0$. There exists $c_{\sigma}>0$ such that:
$\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}].$
The sets $\mathcal{U},\mathcal{Y}$ contain the equilibrium $(r_{y},r_{u})$.
Remark 4.2.
Assumption 4.7 is satisfied by selecting $\hat{u}(N|k)=r_{u}$ and leveraging the predictor’s robustness (Lemma 3.5). Assumption 4.8 is supported by the sensitivity index $I_{s}\leq 0.7$ (Fig. 5), ensuring noise separability and feasibility.
@@ 4.2 Noisy Case: Robustness to Measurement Noise
Under Assumptions 4.1, 4.5, 4.6, 4.7, and 4.8, with a constant admissible equilibrium $(r_{y},r_{u})$, NTDPC is input-to-state stable (ISS) with respect to $\mathbf{\zeta}_{\text{ini}}(k)$. There exist $\beta\in\mathcal{KL}$, $\gamma\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|]\leq\beta(\|\tilde{X}_{\text{ini}}(0)\|,k)+\gamma\left(\sup_{0\leq t<k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}\right).$
@@
Define the non-minimal state as in Definition 4.2:
$\mathbf{z}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}}),\ldots,y(k-1),u(k-T_{\text{ini}}),\ldots,u(k-1)),$
and the error state $\tilde{X}_{\text{ini}}(k)=\mathbf{z}_{\text{ini}}(k)-\text{col}(\mathbf{r}_{y},\mathbf{r}_{u}),$ where $\mathbf{r}_{y}=\text{col}(r_{y},\ldots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\ldots,r_{u})$. The auxiliary function is $W(\mathbf{z}_{\text{ini}}(k))=\sum_{i=1}^{T_{\text{ini}}}l(y(k-i),u(k-i)),$ with stage cost $l(y(i|k),u(i|k))=\|y(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u(i|k)-r_{u}\|_{R_{i}}^{2},$ and $Q_{i},R_{i}\succ 0$. The optimal cost of Problem 4 is $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}.$ Define the stochastic Lyapunov function:

### v2

@@ 4.1 Noise-Free Case
We first analyze the stability of the NTDPC scheme in the ideal noise-free scenario ($\mathbf{\zeta}_{\text{ini}}(k)=0$) while tracking constant admissible references $(r_{y},r_{u})$.
@@
There exist functions $\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ such that for all admissible pairs $(\mathbf{y},\mathbf{u})$:
$\alpha_{1}(\|\operatorname{col}(\mathbf{y}-r_{y},\mathbf{u}-r_{u})\|)\leq l(\mathbf{y},\mathbf{u})\leq\alpha_{2}(\|\operatorname{col}(\mathbf{y}-r_{y},\mathbf{u}-r_{u})\|).$ (47)
@@
For any admissible initial state $\mathbf{z}_{\text{ini}}(k)$, there exists a prediction horizon $N\geq T_{\text{ini}}$ and a terminal control law satisfying a contraction property. Specifically, there exists a function $\rho\in\mathcal{K}_{\infty}$ with $\rho<\operatorname{id}$ such that the terminal cost satisfies:
$l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))\leq(\operatorname{id}-\rho)\circ l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}})).$ (48)
This condition ensures that the cost of the appended terminal state is strictly less than the cost of the oldest data point being discarded from the history window.
@@
The NTDPC optimization (Problem 4) is recursively feasible for all $k\geq T_{\text{ini}}$. In the noise-free case, the optimal slack variable satisfies $\sigma_{y}^{*}(k)=0$. Furthermore, the equilibrium pair $(r_{y},r_{u})$ lies within the interior of the constraint sets $\mathcal{Y}\times\mathcal{U}$.
@@
Under Assumptions 4.1, 4.2, and 4.3, consider the closed-loop system controlled by NTDPC in the absence of noise ($\mathbf{\zeta}_{\text{ini}}(k)=0$). The equilibrium $(r_{y},r_{u})$ is asymptotically stable, implying $\lim_{k\to\infty}\mathbf{y}(k)=r_{y}$ and $\lim_{k\to\infty}\mathbf{u}(k)=r_{u}$.
@@
We construct a Lyapunov candidate using the optimal cost function augmented with a history storage function. Let the optimal cost at time $k$ be $J^{*}(k)$. Since $\sigma_{y}^{*}(k)=0$ (Assumption 4.3), $J^{*}(k)=\sum_{i=0}^{N-1}l(\mathbf{y}^{*}(i|k),\mathbf{u}^{*}(i|k))$.
Define the auxiliary history function $W(\mathbf{z}_{\text{ini}}(k))$ as the accumulated cost of the past $T_{\text{ini}}$ steps:
$W(\mathbf{z}_{\text{ini}}(k))=\sum_{i=1}^{T_{\text{ini}}}l(\mathbf{y}(k-i),\mathbf{u}(k-i)).$ (49)
Consider the Lyapunov function candidate:
$V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k)).$ (50)
Using Assumption 4.1, the history function is bounded below by:
$W(\mathbf{z}_{\text{ini}}(k))\geq\sum_{i=1}^{T_{\text{ini}}}\alpha_{1}(\|\operatorname{col}(\mathbf{y}(k-i)-r_{y},\mathbf{u}(k-i)-r_{u})\|).$ (51)
@@
Since the error state $\tilde{\mathbf{z}}_{\text{ini}}(k)$ is composed of these deviations, there exists $\alpha_{1,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\geq W(\mathbf{z}_{\text{ini}}(k))\geq\alpha_{1,V}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)$. Similarly, using the upper bound $\alpha_{2}$ and the linearity of the predictor $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini}}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)$, the optimal cost $J^{*}(k)$ is bounded by the state magnitude. Thus, there exists $\alpha_{2,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)$.
At time $k+1$, we construct a feasible suboptimal input sequence $\mathbf{u}_{s}(k+1)$ by shifting the optimal sequence from time $k$ and appending the terminal law $\hat{\mathbf{u}}(N|k)$:
$\mathbf{u}_{s}(k+1)=\operatorname{col}(\mathbf{u}^{*}(1|k),\ldots,\mathbf{u}^{*}(N-1|k),\hat{\mathbf{u}}(N|k)).$ (52)
The corresponding cost is $J_{s}(k+1)=J^{*}(k)-l(\mathbf{y}^{*}(0|k),\mathbf{u}^{*}(0|k))+l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))$. By optimality, $J^{*}(k+1)\leq J_{s}(k+1)$.
Now consider the evolution of the history function $W$:
$W(\mathbf{z}_{\text{ini}}(k+1))$ $=\sum_{i=1}^{T_{\text{ini}}}l(\mathbf{y}(k+1-i),\mathbf{u}(k+1-i))$
$=W(\mathbf{z}_{\text{ini}}(k))+l(\mathbf{y}(k),\mathbf{u}(k))-l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}})).$ (53)
Note that in the noise-free prediction, $\mathbf{y}^{*}(0|k)=\mathbf{y}(k)$ and $\mathbf{u}^{*}(0|k)=\mathbf{u}(k)$. Computing the difference $\Delta V(k)=V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))$:
$\Delta V(k)$ $\leq[J^{*}(k)-l(\mathbf{y}(k),\mathbf{u}(k))+l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))]$
$\quad+[W(\mathbf{z}_{\text{ini}}(k))+l(\mathbf{y}(k),\mathbf{u}(k))-l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}}))]$
$\quad-[J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))]$
$=l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))-l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}})).$ (54)
Applying Assumption 4.2 (Terminal Stabilizing Condition):
$\Delta V(k)\leq-\rho\circ l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}})).$ (55)
Since $l(\cdot)$ is positive definite with respect to the deviation from equilibrium, this implies $\Delta V(k)$ is negative definite in terms of the error state. Standard Lyapunov arguments [23] then guarantee that $\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|\to 0$ as $k\to\infty$, confirming asymptotic stability.
We assume that the system is stabilizable and that an admissible input sequence exists that maintains feasibility beyond the prediction horizon.
4.2 Noisy Case: Robustness to Measurement Noise
We now extend the analysis to the case where the measurements are corrupted by noise ($\mathbf{\zeta}_{\text{ini}}(k)\neq 0$). We prove that the NTDPC closed-loop system is Input-to-State Stable (ISS) with respect to the measurement noise $\mathbf{\zeta}_{\text{ini}}(k)$, utilizing the stochastic stability framework for MPC [23, 2].
Assumption 4.5 (Bounded Noise Variance).
The measurement noise is a martingale difference sequence with bounded variance. Specifically, there exists a constant $\bar{c}>0$ such that for all $k\geq 0$:
$\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}\mid\mathcal{F}_{k}]\leq\bar{c},$ (56)
where $\mathcal{F}_{k}$ denotes the filtration generated by measurements up to time $k$.
Assumption 4.6.
[Lipschitz Continuity of the Predictor] The predictor mapping is Lipschitz continuous with respect to the initial condition. There exists a constant $L_{P}>0$ such that for any two initial conditions $\mathbf{z}_{1},\mathbf{z}_{2}$:
$\|\mathbf{P}_{1}(\mathbf{z}_{1}-\mathbf{z}_{2})\|\leq L_{P}\|\mathbf{z}_{1}-\mathbf{z}_{2}\|.$ (57)
Consequently, the deviation in the slack variable due to noise is linearly bounded:
$\|\sigma_{y,s}(k)\|\leq L_{P}\|\mathbf{\zeta}_{\text{ini}}(k)\|,$ (58)
where $\mathbf{z}_{\text{ini},m}(k)=\mathbf{z}_{\text{ini}}(k)+\operatorname{col}(0,\mathbf{\zeta}_{\text{ini}}(k))$.
Assumption 4.7 (Stochastic Terminal Stabilizing Condition).
For any admissible state, there exists a terminal control law $\hat{\mathbf{u}}(N|k)$ and a class $\mathcal{K}_{\infty}$ function $\rho<\operatorname{id}$ such that the expected terminal cost satisfies a contraction:
$\mathbb{E}[l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))\mid\mathcal{F}_{k}]\leq(\operatorname{id}-\rho)\circ\mathbb{E}[l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}}))\mid\mathcal{F}_{k}].$ (59)
Assumption 4.8.
[Recursive Feasibility and Slack Bound] Problem 4 is feasible with probability at least $1-\epsilon$. Furthermore, provided the slack penalty weighting matrix $\Lambda_{y}$ is chosen sufficiently large relative to the noise covariance, the optimal slack variable $\sigma_{y}^{*}(k)$ satisfies a quadratic bound with respect to the noise realization:
[…]
for some constant $c_{\sigma}>0$.
@@ 4.2 Noisy Case: Robustness to Measurement Noise
Assumption 4.7 is typically satisfied by selecting $\hat{\mathbf{u}}(N|k)$ as the steady-state reference $r_{u}$, utilizing the inherent robustness of the predictor established in Lemma 3.4. Assumption 4.8 is empirically supported by the sensitivity index analysis ($I_{s}\leq 0.7$), which ensures sufficient separability between signal and noise subspaces to maintain feasibility.
@@
Under Assumptions 4.1, 4.5, 4.6, 4.7, and 4.8, the closed-loop system generated by NTDPC is Input-to-State Stable (ISS) with respect to the noise sequence $\mathbf{\zeta}_{\text{ini}}(k)$. Specifically, there exist functions $\beta\in\mathcal{KL}$ and $\gamma\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|]\leq\beta(\|\tilde{\mathbf{z}}_{\text{ini}}(0)\|,k)+\gamma\left(\sup_{0\leq t<k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}\right).$ (61)
@@
Consider the stochastic Lyapunov function candidate:
$V(k)=\mathbb{E}[J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))\mid\mathcal{F}_{k}].$ (62)
From the noise-free analysis, $W(\mathbf{z}_{\text{ini}}(k))$ is lower-bounded by $\alpha_{1,W}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)$. The optimal cost $J^{*}(k)$ differs from the noise-free case by the slack term. Using Assumption 4.8 and the boundedness of the noise:
$\mathbb{E}[J^{*}(k)\mid\mathcal{F}_{k}]\leq\alpha_{2,J}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$ (63)
Thus, $V(k)$ satisfies the sandwich bounds:
$\alpha_{1,V}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)\leq V(k)\leq\alpha_{2,V}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$ (64)
We construct the standard suboptimal sequence $\mathbf{u}_{s}(k+1)$ by shifting the optimal input and appending the terminal law. The slack variable for this suboptimal solution, $\sigma_{y,s}(k+1)$, absorbs the prediction error due to noise. From Assumption 4.6, its expected cost is bounded:
$\mathbb{E}[\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}\mid\mathcal{F}_{k}]\leq\lambda_{\max}(\Lambda_{y})L_{P}^{2}\bar{c}.$ (65)
Define $\bar{C}_{\text{noise}}=(c_{\sigma}+\lambda_{\max}(\Lambda_{y})L_{P}^{2})\bar{c}$. The expected difference in the Lyapunov function is:
$\mathbb{E}[V(k+1)-V(k)\mid\mathcal{F}_{k}]$ $\leq\mathbb{E}\left[l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))-l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}}))\right$.
$\quad\left.+l(\mathbf{y}(k),\mathbf{u}(k))-l(\mathbf{y}^{*}(0|k),\mathbf{u}^{*}(0|k))\mid\mathcal{F}_{k}\right]+\bar{C}_{\text{noise}}.$ (66)
The first line is bounded by $\rho(\cdot)$ via Assumption 4.7. The crucial term is the difference in the second line: $\Delta_{\text{stage}}=l(\mathbf{y}(k),\mathbf{u}(k))-l(\mathbf{y}^{*}(0|k),\mathbf{u}^{*}(0|k))$. In the noise-free case, $\mathbf{y}(k)=\mathbf{y}^{*}(0|k)$, so this term vanishes. In the noisy case, the measured output $\mathbf{y}(k)$ differs from the predicted output $\mathbf{y}^{*}(0|k)$ due to the noise realization at time $k$. Since $l(\cdot)$ is locally Lipschitz with constant $L_{l}$ on the compact constraint set: